OpenStudy (heyitsme541):

The first term of an arithmetic progression is -3 and the common difference is 8. What is the 28th term?

jimthompson5910 (jim_thompson5910):

We're told that `The first term of an arithmetic progression is -3` so this means \(\Large a_1 = -3\) Also, we're given ` the common difference is 8` which tells us that \(\Large d = 8\) Do you agree with everything so far?

OpenStudy (heyitsme541):

Yes

jimthompson5910 (jim_thompson5910):

what we do next is plug those values into the formula below and simplify \[\Large a_n = a_1 + d(n-1)\] \[\Large a_n = -3 + 8(n-1)\] \[\Large a_n = -3 + 8n-8\] \[\Large a_n = 8n-11\] this is our nth term formula

jimthompson5910 (jim_thompson5910):

It asks: `What is the 28th term?` so they want us to plug in \(\Large n = 28\) into the nth term formula \[\Large a_n = 8n-11\] \[\Large a_{28} = 8(28)-11\] \[\Large a_{28} = ???\] I'll let you finish up

OpenStudy (heyitsme541):

So it'd be 213? Thanks!

jimthompson5910 (jim_thompson5910):

`So it'd be 213? ` yes, correct